Can you please help me with this: Sam has two rectangle rooms in her house, the combined floor area of booth rooms is 321 square ft, the first room is 13 ft long and 12 ft wide, second room is 11 ft wide. What is the length in feet of the second room?
Answer by mananth(16946)

One room area = 13*12 = 156 square ft

Let length of the other room be x
Other room area = 11 *x
11x+156=321
11x=321-156
11x=165
/11
x=15 feet